a-2b/a+2b = 1k+1
solve for a:

Answer by monika_p(71)

a-2b/a+2b = 1k+1 multiply both sides by (a+2b)
a-2b=(k+1)(a+2b) multiply two () on the right side
a-2b=ak+2bk+a+2b shift terms with a on the left side and all rest on the left
a-ak-a=2bk+2b+2b
-ak=2bk+4b divide both sides by -k
a=-2b(k+2)/k